OpenStudy (anonymous):

Find the area of the vegan bounded by the parabola y=x^2 the tangent line to this parabola at (1,1) and the x-axis.

OpenStudy (anonymous):

area of the vegan? :) does your parabola eat no animal products?

OpenStudy (anonymous):

oops i meant area of the region bounded by the parabola

OpenStudy (anonymous):

Sorry Shyrene, idk. But since a parabola is a type of line, won't it be infinite, therefore haveng an infinite area?

OpenStudy (anonymous):

idk? that's why im here asking for some help?.... if only they actually had examples like these in the book.

myininaya (myininaya):

so we need to find the tangent line at (1,1) of y=x^2 do you know how to do that?

myininaya (myininaya):

let me know what you get for the tangent line at (1,1) of y=x^2 ok?

myininaya (myininaya):

and i will tell you if it is right or not. and did you see my drawing above?

OpenStudy (anonymous):

tangent line y=2x?

myininaya (myininaya):

you got y'=2 at x=1. so we have the slope of y=mx+b. so we have y=2x+b. we need to find b the y-intercept. So we know a point (1,1) on the line. so we have 1=2(1)+b. We solve this equation for b and get b=-1. so the tangent line to x^2 at (1,1) is y=2x-1

OpenStudy (anonymous):

thank you i actually understood that.

myininaya (myininaya):

np